OpenStudy (anonymous):

Find dy/dx using log differentiation. y = x^(x-1)

OpenStudy (aum):

\[ y = x^{x-1} \\ \log(y) = \log(x)^{x-1} = (x-1)\log(x) \\ \text{Implicit Differentiation on both sides:} \\ \frac{dy}{y} = ((x-1)\frac 1x + \log(x))dx\\ \frac{dy}{dx} = y * ((x-1)\frac 1x + \log(x))\\ \frac{dy}{dx} = x^{x-1} * ((x-1)\frac 1x + \log(x))\\ \]

OpenStudy (aum):

\[ \frac{dy}{dx} = x^{x-1} * ((x-1)\frac 1x + \log(x))\\ \frac{dy}{dx} = x^{x-1} * ((1-\frac 1x) + \log(x))\\ \frac{dy}{dx} = x^{x-1} - x^{x-2} + x^{x-1} \log(x)\\ \]

OpenStudy (anonymous):

Thank you!